\begin{thebibliography}{99}

\bibitem{CZ}
Q.-M. Cheng and L. Zeng, \newblock{Eigenvalues of the Witten-Laplacian on compact Riemannian manifolds.} \newblock{\em arXiv: math. DG/1304.3212v1.}

\bibitem{EI}
A. El~Soufi and S. Ilias, \newblock{Second eigenvalue of Schr\"{o}dinger operators and mean curvature.} \newblock{\em Comm. Math. Phys.,} 208:  761--770, 2000.

\bibitem{EHI}
A. El~Soufi, E.~M. Harrell, II, and S. Ilias, \newblock{Universal
  inequalities for the eigenvalues of {L}aplace and {S}chr\"odinger operators
  on submanifolds.} \newblock{\em Trans. Amer. Math. Soc.,} 361 (5): 
  2337--2350, 2009. 

\bibitem{IM1}
S. Ilias and O. Makhoul, \newblock{"Universal" inequalities for the eigenvalues of the Hodge de Rham Laplacian.}  \newblock{\em Ann. Global Anal. Geom.,} 36 (2): 191--204, 2009. 

\bibitem{IM2}
S. Ilias and O. Makhoul, \newblock{A generalization of a Levitin and Parnovski Universal inequality for eigenvalues.} \newblock{\em J. Geom. Anal.,} 22:  206--222, 2012. 

\bibitem{S}
A. G. Setti, \newblock{Eigenvalue estimates for the weighted Laplacian on a Riemannian manifold.} \newblock{\em Rend. Sem. Mat. Univ. Padova.,} 100:  27--55, 1998.

\bibitem{WX}
Q. Wang and C. Xia, \newblock{Universal bounds for eigenvalues of Schr\"{o}dinger operator on Riemannian manifolds.} \newblock{\em Ann. Acad. Scient. fennicae math.,} 33: 319--336, 2008.

\end{thebibliography}

\documentclass[11pt,a4paper,reqno]{amsart}
\usepackage{amsmath}
\usepackage{amssymb,latexsym}
\usepackage[dvips]{graphicx}
\usepackage{color}

\begin{document}
\medskip
  \title{ Remark on the paper "On the eigenvalues of the Witten-Laplacian on compact Riemannian manifolds " by Q.-M. Cheng and L. Zeng}
\author{  S. Ilias}
\date{13/04/13}

\maketitle
\indent 
 On April 11, 2013, a preprint was posted on arXiv by Q.-M. Cheng and L. Zeng \cite{CZ}, about eigenvalues of the Witten-Laplacian acting on functions. However, it is very well known that the Witten Laplacian $$\Delta_{f}:=\Delta+<\nabla f,.>$$ on a compact measure metric space $(M,g,e^{-f}\,dv )$ is unitarily equivalent to the Schr\"{o}dinger operator $\Delta+\frac{1}{2}\Delta f+\frac{1}{4}|\nabla f|^{2}$ on $(M,g)$ and it has thus the same spectrum (see for instance \cite{S}). Therefore, all the results of this above mentioned preprint are immediate consequences of that concerning the Schr\"{o}dinger operator  $\tilde{\Delta}_{f}:=\Delta+\frac{1}{2}\Delta f+\frac{1}{4}|\nabla f|^{2}$ on $(M,g)$ already obtained in the papers \cite{EI, EHI, IM1, IM2, WX},  provided replacing the potential $q$ or $V$ in that papers by $\frac{1}{2}\Delta f+\frac{1}{4}|\nabla f|^{2}$   (obviousely, this is in particular, the case for the self shrinkers).\\

\indent Using the same observation of unitary equivalence above, we also note that a lot of published results concerning eigenvalues of the Laplace-Witten operator, can be deduced from that concerning Schr\"{o}dinger operators.
 
\textsc{Said Ilias}

{\em Universit\'e F. Rabelais, D\'ep. de Math\'ematiques, Tours

ilias@univ-tours.fr}

\end{document}